OpenStudy (anonymous):

240 oranges were shared between Dave, Karen, and Sam. In the ratio 4:10:6. How many Oranges did Dave receive if Sam got 72?

OpenStudy (ohohaye):

Try dividing 72 by 6, then multiply that number by 4 and you will get David's portion of the oranges

OpenStudy (anonymous):

This is where I belong

OpenStudy (anonymous):

Hi Ohohaye. Thanks for your response.The answer I got is 48. Thanks for your help! Could you help me with a couple more? I have answers for two.

OpenStudy (anonymous):

Add the total ratio count. 4+10+6=20 240/20=12 Sam=12*6=72 Dave=12*4=?

OpenStudy (anonymous):

I can help you with the other questions if you start them new.

OpenStudy (anonymous):

Thanks OrangeMaster! 1750 shared between 3 people. Ratio 3:5:8. What is the amount of the smallest portion. A population of 120 students is in the ratio of 1:9/ How many students are there to 4 teachers?